### v1

# Investigation of the Impact of Synthetic Training Data in the Industrial Application of Terminal Strip Object Detection

Abstract
In industrial manufacturing, numerous tasks of visually inspecting or detecting specific objects exist that are currently performed manually or by classical image processing methods. Therefore, introducing recent deep learning models to industrial environments holds the potential to increase productivity and enable new applications. However, gathering and labeling sufficient data is often intractable, complicating the implementation of such projects. Hence, image synthesis methods are commonly used to generate synthetic training data from 3D models and annotate them automatically, although it results in a sim-to-real domain gap. In this paper, we investigate the sim-to-real generalization performance of standard object detectors on the complex industrial application of terminal strip object detection. Combining domain randomization and domain knowledge, we created an image synthesis pipeline for automatically generating the training data. Moreover, we manually annotated 300 real images of terminal strips for the evaluation. The results show the cruciality of the objects of interest to have the same scale in either domain. Nevertheless, under optimized scaling conditions, the sim-to-real performance difference in mean average precision amounts to 2.69% for RetinaNet and 0.98% for Faster R-CNN, qualifying this approach for industrial requirements.
1 INTRODUCTION
Currently, various research fields in machine learning (ML) are evolving rapidly, setting new benchmarks and demonstrating the potential benefit of this technology for society. In particular, the availability of computing power and large-scale public datasets advanced the development and initiated the era of deep learning (DL). However, the amount of data required to train DL models limits the transferability of current state-of-the-art approaches to real-world applications in which data is scarce and difficult to gather [Tsirikoglou et al., 2020], such as computer vision (CV) applications in industrial manufacturing.
The main challenge arises from the manual effort to collect and label the training images for supervised learning. For instance, the creators of Microsoft Common Objects in Context (MS COCO), a commonly used benchmark dataset for object detection and semantic segmentation, stated in [Lin et al., 2014] that it took around 60,000 worker hours to gather and annotate these images, which is intractable for most industrial applications. Moreover, industrial settings are dynamic, so the environment and the objects of interest may change, forcing the training data to be dynamically adaptable within a reasonable time. Therefore, recent works focused on the usage of simulated images for model training and investigated their impact on model performance in real-world test cases [Akar et al., 2022, Ljungqvist et al., 2023, Zhu et al., 2023].
For industrial manufacturers, so-called synthetic training data offers the following advantages:
  • Accurate 3D models for image synthesis are often available from computer-aided design (CAD), used during the construction of products.
  • Theoretically, an arbitrary amount of training images can be generated using an automated image synthesis pipeline.
  • Ground-truth annotations can be calculated automatically, eliminating the human source of error and saving costs for manual annotation.
  • The training data can be generated without real-world biases by adapting parameters in the image synthesis pipeline.
However, the domain gap between synthetic and real images often causes the DL models to perform worse in real-world test cases, so numerous approaches have been proposed to close this gap.
Apart from the usual data augmentation techniques and the related cut-and-paste approach [Dvornik et al., 2018], domain adaptation and domain randomization (DR) are the two main concepts currently used in this area based on rendered images of 3D models. The former maps one domain into the other or both into an intermediate third domain, using, e.g., GANs [Rojtberg et al., 2020, Shrivastava et al., 2017, Wang et al., 2022]. Thus, domain adaptation still relies on real images, which might be unsuitable for industrial use cases. In contrast, DR is based on random, even unrealistic variations, e.g., of textures, backgrounds, or lighting conditions, to suggest to the DL model that real images are just another variation of the learned domain [Tobin et al., 2017]. Hence, it aims to generalize directly to the real world and enables fully synthetic training, which has already shown promising results in different use cases, such as detecting cars [Tremblay et al., 2018], small load carriers [Mayershofer et al., 2021], and turbine blades at a manual working station [Eversberg and Lambrecht, 2021]. Nevertheless, experiments in industrial environments were conducted under highly simplified conditions, including a small amount of well-separable classes and considering only similar-sized objects. The expectable performance of state-of-the-art approaches on more complex industrial tasks remains unclear.
In this paper, we assess the particularly challenging object detection of terminal strips using 3D models by Phoenix Contact that are publicly available on the company’s homepage [^1] https://www.phoenixcontact.com/en-pc/products/terminal-blocks/feed-through-terminal-blocks-multi-conductor-terminal-blocks-and-multi-level-terminal-blocks. The investigation is motivated by two applications: identification of already mounted terminal blocks in electrical cabinets and quality control in terminal strip assemblies. The first application is relevant, as the product information is printed on the side of the terminal blocks and occluded by neighboring components in a mounted state. Thus, experts must identify certain terminal blocks to avoid demounting. The second application of quality control stems from the fact that the product portfolio for terminal strips is manifold, including over 10,000 components. This causes manual work steps during terminal strip assembly. Since manual processes are prone to human error and experts are not always available, a ML-based detection system would prevent errors and save time in both cases.
We approached this task by building an automated image synthesis pipeline and generating 30,000 synthetic images of random terminal strips. Moreover, we annotated 300 handmade real images for evaluation and provide baseline results of two standard object detectors, entirely trained on synthetic images and tested on real images. As we found that scaling is crucial in this specific use case, we also show how synthetic data can be used to learn a suitable image preprocessing that significantly improves the detection performance. To motivate further developments on this use case, we made our dataset publicly available on Zenodo [^2] https://zenodo.org/records/10674995 for research purposes.
The remainder of this paper is structured as follows. Section 2 points out usual image synthesis methods and available synthetic image datasets to show recent developments in sim-to-real DL. Then, we describe in Section 3 how DR and domain knowledge are combined in our pipeline to create a synthetic image dataset for terminal strip object detection. While we detail our experiments to assess the sim-to-real domain gap and present the corresponding results in Section 4, Section 5 discusses the significance of the results for industrial manufacturers. The paper is concluded in Section 6 also giving a brief outlook for future work.
2 RELATED WORK
This section briefly discusses popular synthetic image datasets and recent image synthesis methods to cover the current state-of-the-art of using simulated data for visual DL.
2.1 Synthetic Image Datasets
Although annotated synthetic image datasets can be generated much faster than their real-world counterparts, they are still outnumbered and only available for specific use cases. One very active research area dealing with simulated training data is autonomous driving, in which different approaches are used for dataset creation. In [Richter et al., 2016], e.g., the authors exploited the video game engine of GTA5 to obtain 25,000 photorealistic street scene images with pixel-level segmentation masks. By substituting $\nicefrac{{2}}{{3}}$ of the real training images, they could improve the performance of a semantic segmentation model, which demonstrates the potential benefit of synthetic data.
Contrarily, the creators of Virtual KITTI semi-automatically reconstructed real-world street scenes from the original KITTI dataset [Geiger et al., 2012] and varied several parameters, such as the weather conditions, resulting in 35 synthetic videos with approximately 17,000 frames labeled for various video understanding tasks [Gaidon et al., 2016]. Their results show that a combination of synthetic pre-training and real fine-tuning yields better performance than each separate approach in multi-object tracking, supporting the importance of simulated data.
Another notable street scene dataset is Synscapes [Wrenninge and Unger, 2018] which represents the synthetic counterpart to Cityscapes [Cordts et al., 2016]. In contrast to the previously mentioned datasets, Synscapes was created using a parameter-based random generator to build a new scene for each image. Furthermore, the developers leveraged photorealistic rendering techniques and found that high image realism is necessary to bridge the sim-to-real domain gap [Wrenninge and Unger, 2018]. Nevertheless, the application of autonomous driving is out of scope for industrial manufacturers, so the observed results may not apply to their use cases.
The amount of synthetic image datasets for industrial requirements is small, even though several approaches for image synthesis have been proposed and evaluated in small-scale industrial test cases, as discussed in the following subsection. Two recently published datasets are SIP-17 [Zhu et al., 2023] and SORDI [Akar et al., 2022], aiming to close this gap. SIP-17 is a synthetic dataset for industrial parts classification consisting of 17 objects in six different use cases to allow research in close-to-reality scenarios. The experiments reveal that classification is difficult when classes share a similar albedo or have a comparable shape [Zhu et al., 2023], which is typical for industrial objects.
While SIP-17 contains 33,000 images, SORDI was presented as the first large-scale synthetic image dataset for the industry with 200,000 images at the publication date. It is characterized by eight industrial assets in 32 fully synthetic factory scenarios and enables the sim-to-real benchmarking of vision models for intelligent robots in factory environments [Akar et al., 2022]. However, SORDI cannot serve the entire range of industrial applications, so further work is required to cover, e.g., quality control in assembly and object tracking.
To conclude, it is difficult for industrial manufacturers to estimate how a synthetically trained DL model would perform in a specific use case, as there is not enough data available to benchmark sim-to-real methods in different industry-related applications. Thus, we provide our dataset to enable further investigations in this area with a focus on narrow industrial objects that are densely arranged and partly occlude each other.
2.2 Image Synthesis for Deep Learning
Generating synthetic training images for supervised learning tasks requires suitable design choices to improve the sim-to-real generalization performance. These choices include textures of the objects of interest, the image background, different viewpoints, lighting conditions, etc., potentially making the modeling of virtual scenes very time-consuming and complex as well. Especially when the reality should be closely approximated, modeling the entire range of possible real-world variations might eliminate the advantage of saving time for data annotation. Hence, there are various approaches to apply image synthesis that attempt to minimize the domain gap while avoiding excessive modeling effort. For instance, the authors of [Dwibedi et al., 2017] used segmentation masks of real images to cut out the objects of interest and place them onto random background images, which can be seen as image synthesis. This idea was later improved in [Dvornik et al., 2018] proposing a context model that provides patches of the background images likely to contain certain objects to create more realistic training data. Although this method is very efficient and showed promising results in object detection [Dvornik et al., 2018], it relies on real-world data which might not be available or too expensive to gather in industrial applications. Alternatively, the appearance of real-world objects can be approximated utilizing accurate 3D models and modern rendering techniques such as in [Su et al., 2015]. Here, the authors overlaid background images with object renderings, similar to the cut-and-paste approach but completely removing the necessity of annotated real-world data.
Since real images are highly affected by the environment in which they are taken, the appearance of objects changes with their surroundings, impeding model prediction. Hence, a common approach is to randomize the environment of the objects of interest during image synthesis to train the DL model to deal with these variations. However, the authors of [Tobin et al., 2017] hypothesized that complete DR, including even unrealistic variations of the objects of interest, helps to close the sim-to-real domain gap as the DL model is forced to focus on the object’s shape rather than its color or glossiness. Therefore, they also randomized the textures of the 3D models and obtained promising results in the localization of simple geometric objects, enabling a robot to grab these objects in a real-world test [Tobin et al., 2017].
Based on this idea, several experiments have been conducted, assessing different degrees of DR and how domain knowledge can be used during simulation to improve real-world performance. For instance, [Prakash et al., 2019] introduced context splines to the synthetic scenes to ensure that cars are always positioned on streets and [Mayershofer et al., 2021] proposed to define object relations via configuration files to set the objects of interest into more realistic contexts. However, these methods were evaluated on small individual test cases, so the results are hardly comparable and also partly contradicting. This was addressed by the authors of [Eversberg and Lambrecht, 2021] who compared numerous DR approaches in the application of detecting turbine blades at manual working stations to provide a guideline on designing image synthesis pipelines for industrial use cases. They observed it to be most beneficial to focus on realistic lighting conditions and object textures while the image background and distractors can be randomized completely [Eversberg and Lambrecht, 2021].
Complementarily, in our research, we follow these guidelines and make a few application-specific changes to demonstrate the sim-to-real performance of state-of-the-art image synthesis techniques in the challenging industrial application of terminal strip object detection.
3 IMAGE SYNTHESIS PIPELINE
The image synthesis pipeline takes 3D models and outputs annotated synthetic images of terminal strips, exploiting domain knowledge to set the objects of interest into a more realistic context. In this application, the most general characteristics we focus on are the following:
  • Terminal strips consist of several terminal blocks mounted next to each other on a DIN rail.
  • The terminal blocks are equipped with additional accessories such as plug-in bridges, test adapters, markings and end covers.
  • Neighboring terminal blocks often are of the same type, resulting in a group-like structure.
  • Terminal strips are mounted on flat surfaces such as back panels of electrical cabinets.
  • The most informative view of a terminal strip is the front view, as illustrated in Figure 1.
Since it has been shown to be beneficial to include domain knowledge in DR [Prakash et al., 2019, Mayershofer et al., 2021], we use this information in our pipeline along with standard randomization techniques to obtain random synthetic images that are contextually close to reality. In this section, we provide details about the different generation steps we implemented using the open-source rendering software Blender [^3] Blender 3.3: https://www.blender.org.
Figure 1: Example rendering of a terminal strip randomly generated by the process described in Subsection 3.1. It shows the centered front view that is most informative in this use case.
3.1 Creation of Terminal Strips
In this paper, we consider 36 terminal blocks, including two connection technologies, three wire sizes, and two to four connection points to cover almost identical and easily distinguishable objects. Furthermore, the terminal blocks are randomly equipped with plug-in bridges, test adapters, markings, and end covers, and different realistic textures are used. Therefore, we manually created virtual materials by adapting roughness, specular, metallic, and RGB values in Blender until the renderings looked similar to real reference images.
The first step for creating a terminal strip regarding domain knowledge is to place the 3D model of a DIN rail in the center of the virtual scene. After that, uniformly sampled terminal blocks are virtually mounted next to each other on the DIN rail until the number of objects exceeds a threshold that restricts the size of the terminal strip. Each selected 3D model of a terminal block is thereby inserted up to five times to model the group-like structure of terminal strips, and the created virtual materials are randomly assigned. Since gray housings are more common in reality than those of other colors, we set the probability for the gray virtual material to 80%, while the remaining colors uniformly share 20%. Moreover, most real-world terminal blocks are one-sided open, requiring an end cover if the neighboring terminal block does not entirely cover the open side to ensure electrical insulation. This is implemented in the pipeline by inserting an end cover every time the height or depth of the selected terminal block is less than 90% of the previous one.
Apart from wiring, the main distractors of terminal strips are their additional accessories, such as plug-in bridges, test adapters, and markings. Although the authors of [Eversberg and Lambrecht, 2021] found that realistic distractors do not necessarily improve the performance of DL models, we decided to stay with realism for two reasons. First, the modeling effort is almost identical to simple, unrealistic distractors, as the required 3D models are available, and the virtual materials are similar to the materials used for the terminal blocks. Second, some applications may focus on detecting these accessories, e.g., to determine whether plug-in bridges are plugged correctly. Thus, each terminal strip is randomly equipped with markings, plug-in bridges, and test adapters according to its configuration.
Starting with the markings, we decide for every marking point individually whether the 3D model of the corresponding marking is inserted by sampling from a Bernoulli distribution with a probability of success of 70%. In contrast, randomly placing plug-in bridges is more difficult since not all types of terminal blocks can be bridged, and the length of plug-in bridges is variable. Hence, we iterate over all shafts and determine the bridgeable ones to sample realistic bridge positions and different lengths afterward. Furthermore, red and blue plug-in bridges are considered with a probability of 70% for red, as this color appears more often in reality. Regarding test adapters, the 3D models are randomly placed in the remaining shafts with a probability of 10%, varying between four colors, and 20% of the terminal strips are rendered without test adapters.
Finally, terminal strips are often enclosed by so-called end clamps, so the pipeline is implemented to generate 50% of the terminal strips containing a starting and an ending 3D model of an end clamp. Figure 1 shows a sample rendering of a terminal strip generated by the described randomization procedure.
3.2 Viewpoints
Considering different viewpoints in object detection and related ML-based CV tasks is essential, as the objects of interest usually appear arbitrarily rotated and scaled in real-world images. However, domain knowledge often enables the exclusion of unrealistic settings during image synthesis, reducing the number of images required to cover frequently occurring viewpoints. As mentioned previously, the most informative view of terminal strips is their front view, illustrated in Figure 1, which also is the only available view when the terminal strip is already mounted on a flat surface. Although in a production environment, such as in quality control, a camera could be fixed in a well-centered position, slightly varying viewpoints occur when an image is taken by hand, e.g., in front of an electrical cabinet, which is the more general case we focus on.
In our pipeline, we place a virtual camera in Blender perfectly centered in front of the generated terminal strip and directed to its center to start with the optimal front view. Then, we independently add 3D Gaussian noise to the camera’s position and orientation, resulting in a majority of viewpoints close to the reference view while reasonably deviating viewpoints are also covered. As a side effect, this approach captures slightly different scales of the objects of interest, as the distance between the camera and the terminal strips varies. To show the outcome of this viewpoint randomization, Figure 2 visualizes 1,000 sampled camera positions in the form of a point cloud.
Figure 2: Visualization of 1000 sampled camera positions in the form of a point cloud to illustrate the procedure of randomizing viewpoints.
3.3 Lighting Conditions
In recent image synthesis methods for visual DL applications, there are two common approaches to randomize the lighting conditions: using parameterizable, virtual light sources and implementing image-based lighting. The former is very flexible and allows accurate modeling of the desired illumination. However, creating real-world conditions is laborious since light rays of different colors and intensities might enter the scene from various directions, and reflections should also be considered. In contrast, image-based lighting maps a $360^{\circ}$ image on a globe and uses the pixel values to illuminate the 3D models as if they were in the captured environment. This is less customizable but reduces the creation of complex lighting conditions to the selection of suitable images and has shown to be superior when only little effort is invested in parameterizing virtual light sources [Eversberg and Lambrecht, 2021].
To keep the modeling effort small, we selected 46 high dynamic range images (HDRIs) from Poly Haven [^4] https://polyhaven.com/hdris representing $360^{\circ}$ indoor scenes with an industrial context or medium contrast, as shown in Figure 3. During dataset creation, a new HDRI is uniformly sampled for each generated terminal strip and imported such that the back side of the DIN rail faces the ground of the captured environment. This is reasonable since it excludes the presence of light sources behind the DIN rail corresponding to the real-world conditions when the terminal strip is mounted, e.g., in an electrical cabinet. To further increase the variety of possible illuminations, the HDRIs are randomly rotated around the vertical axis with a resolution of $1^{\circ}$ so that $46\cdot 360=16,560$ at least slightly different lighting conditions can occur in the resulting dataset.
Figure 3: Example $360^{\circ}$ HDRI from Poly Haven used for the image-based lighting.
3.4 Image Background
Similar to the lighting conditions, synthetic image backgrounds can be virtually created and rendered along with the objects of interest [Akar et al., 2022] or they rely on real images [Eversberg and Lambrecht, 2021]. On the one hand, modeling the entire background leads to more realistic images, as all objects can be related, and visualizing realistic shadows is possible. On the other hand, random image-based backgrounds have shown promising results [Eversberg and Lambrecht, 2021] and are less complex, which is more suitable for industrial manufacturers. Nevertheless, the authors of [Alghonaim and Johns, 2021] found that realistic shadows significantly improve the performance of DL models in a 6D pose estimation task. Therefore, we decided to follow a compromising strategy, reusing the HDRIs of the environment lighting and visualizing the shadow of the terminal strip as if it were mounted on a flat surface.
Figure 4: Illustration of the semi-ellipsoid used to visualize the shadows of the terminal strips and onto which the HDRIs are mapped to create the image background.
The implementation is based on a surrounding semi-ellipsoid, as illustrated in Figure 4, that serves as a shadow catcher. It is positioned so that the flat side represents the mounting surface for the DIN rail to which the terminal strip is centered, allowing the objects of interest to cast shadows on the background. To insert the HDRI, the same image from the environment lighting is mapped onto the surface of the semi-ellipsoid, and the respective rotation around the vertical axis is applied. As a result, the background of a synthetic image consists of the ground of the environment captured by the HDRI with additional shadows cast by the terminal strip, which is more realistic than using fully randomized backgrounds and less expensive than modeling the entire scene. See the appendix for synthetic sample images generated using this approach.
3.5 Annotations
The main advantage of generating synthetic images for supervised learning is that ground-truth annotations can be calculated automatically and almost for free. In object detection tasks, these annotations consist of a class label and the position of the object, represented as an axis-aligned bounding box. However, even in synthetic environments, accurate bounding boxes can be hard to provide in so-called crowded scenes, meaning that objects are densely arranged and occlude each other. This also applies to terminal strips, as can be seen in Figure 5.
While scenario (a) shows an unproblematic case of a well-centered front view of two terminal blocks, image (b) demonstrates that a slight camera rotation of $5^{\circ}$ around the viewing direction already widens the bounding boxes, such that they include half of the neighboring terminal blocks. Moreover, scenarios (c) and (d) illustrate that components can occlude each other, and terminal blocks with a roof-like shape are hard to annotate, depending on the camera position.
Figure 5: Four different annotation scenarios that illustrate the difficulty of determining accurate bounding boxes for narrow objects, even in image synthesis.
The comparison shows that including all visible parts of a terminal block into its bounding box inevitably leads to highly overlapping annotations, which is critical for most DL-based detection models. The major problem arises from the non-maximum suppression (NMS) algorithm commonly used in object detection. It is a post-processing step that iteratively selects the detection with the highest prediction score and removes all highly overlapping detections with the same predicted class label based on a threshold. In crowded scenes, NMS often results in a large number of false negatives, i.e., undetected objects, which compromises the overall detection performance [Jiang et al., 2023]. Therefore, we decided to restrict the overlaps of terminal block bounding boxes to 20% of the width of the thinner component, encouraging the DL models to learn tight positional predictions with reasonable overlaps during training. Additionally, we provide annotations for the end clamps, end covers, plug-in bridges, and test adapters to enable applications in which accessories should be detected.
4 EVALUATION
In this section, we evaluate the sim-to-real performance of standard object detection models in the industrial application of terminal strip object detection using the proposed image synthesis pipeline. Therefore, we selected the two standard model architectures, RetinaNet [Lin et al., 2017] and Faster R-CNN [Ren et al., 2017], representatively for the common single-stage and two-stage detectors and used different backbones according to the TensorFlow Object Detection (TFOD) framework [Huang et al., 2017]. As we found RetinaNet with a ResNet50 [He et al., 2016] backbone and Faster R-CNN with an Inception-ResNet [Szegedy et al., 2017] backbone to perform best in previous experiments without accessories, we leveraged these particular DL models for our present investigation of the domain gap.
The evaluation is based on the mean average precision (mAP) metrics mAP@0.5 and mAP@[0.5:0.95], as defined for the MS COCO [Lin et al., 2014] challenge. While the first considers a single intersection over union (IoU) threshold of 0.5, the second averages over ten equidistant IoU thresholds ranging from 0.5 to 0.95 to evaluate the predicted position of the objects of interest more strictly. Thus, it can be distinguished between an overall poor detection performance and correct detections with slightly inaccurate bounding boxes.
4.1 Datasets
For our experiments, we generated a synthetic dataset of 30,000 randomly configured terminal strip images at a resolution of $1024\times 512$ pixels using the image synthesis pipeline. It contains 36 frequently occurring terminal blocks in five different colors and suitable plug-in bridges, test adapters, end covers, and end clamps as additional accessories. The markings are only considered as distractors and not part of the detection, so 40 detection classes, composed of very similar and well-separable objects, are provided in total. The average number of objects per image is 25, whereas the accessories are overrepresented compared to the evenly distributed terminal blocks, as they occur in almost every image. The dataset is split into 79% train, 1% evaluation, and 20% test images to enable testing on synthetic images required to compare the performances of the DL models in simulation and reality.
To perform the real-world test, we used the first part of the image synthesis pipeline, as described in Subsection 3.1, to obtain 100 parts lists of random terminal strips and build them in reality. After that, we took three pictures of each terminal strip from different viewpoints and manually labeled the resulting 300 images for object detection. Hence, the image contents in the synthetic and the real-world datasets are very similar, enabling a meaningful investigation of the sim-to-real domain gap. Since some terminal blocks only marginally differ in their width, e.g., UT 4 and UT 6, we expect the DL models to confuse these classes in either domain.
4.2 Experiments
In our experiments, we leveraged the implementation of RetinaNet and Faster R-CNN available in the TensorFlow 2 Detection Zoo [^5] https://github.com/tensorflow/models/blob/master/research/object˙detection/g3doc/tf2˙detection˙zoo.md of TFOD. In this framework, the models can be configured via configuration files and fine-tuning from ImageNet [Deng et al., 2009] or MS COCO [Lin et al., 2014] pre-training is possible. Since we aim to provide the performance of standard DL models, we kept the default configuration except from the following changes:
  • The target size of the preprocessing image resizer was adapted to the size of the synthetic images.
  • All data augmentation methods were removed since the dataset could be extended by generating more synthetic images if necessary, and further variations are also realizable in the image synthesis pipeline.
  • The aspect ratios of the anchor boxes were adapted to the most frequent bounding box ratios in the synthetic training dataset.
  • The number of training steps was set to 50,000.
  • The batch size for RetinaNet was set to eight and for Faster R-CNN to four due to hardware limitations.
  • ImageNet pre-training was selected.
Using these configurations, the RetinaNet and Faster R-CNN detection models were trained on the synthetic training data and evaluated on the synthetic and real test images to quantify the sim-to-real domain gap.
As we found the real images to show the terminal strips on a smaller scale than the synthetic training images, we further experimented with rescaling the real-world data. Therefore, we considered three different approaches to determine the rescaling factors. First, a constant factor of 1.5 was applied to all real images, as it seemed to scale the majority into the range of the synthetic images. Second, the image synthesis pipeline was used to generate 1,000 terminal strip images of various smaller scales annotated with the corresponding rescaling factors to fine-tune a simple ResNet50 regression model from ImageNet pre-training and predict the scaling factor for each real image separately. Finally, we considered the trained DL models as black-box functions and applied a Bayesian optimization algorithm to the real test images to obtain the best detection performance of RetinaNet and Faster R-CNN depending on the scaling factors. Accordingly, we refer to these experiments as real (1.5), real (ResNet), and real (Opt) in the following. Note that the latter approach is not applicable in a real-world application, as it requires ground-truth annotations. However, it demonstrates the potential performance when accurate predictions for image resizing are available and allows the evaluation of the ResNet50 regression model.
4.3 Results
The performances of RetinaNet and Faster R-CNN in our experiments are listed in Tables 1 and 2, respectively. As you can see, both models reach 99% $mAP@0.5$ and more than 92% $mAP@\left[0.50:0.95\right]$ on the synthetic test images, showing that terminal strip object detection is well-solvable in the simulated domain, and even very similar terminal blocks can be distinguished. Considering this as the reference performance, the sim-to-real domain gap seems to be huge in the case of unscaled real test images since the better Faster R-CNN only achieves 73.80% $mAP@0.5$, and the performance of both models drops by more than 20% when the predicted bounding boxes are evaluated more strictly.
Table 1: Comparison of the sim-to-real detection performances of RetinaNet based on different scaling factors.
Test dataset $mAP@0.5$ $mAP@\left[0.50:0.95\right]$
synthetic $99.26\%$ $92.57\%$
[…]
real (1.5) $70.02\%$ $46.42\%$
real (ResNet) $90.75\%$ $63.26\%$
real (Opt) $96.57\%$ $68.23\%$
Table 2: Comparison of the sim-to-real detection performances of Faster R-CNN based on different scaling factors.
Test dataset $mAP@0.5$ $mAP@\left[0.50:0.95\right]$
synthetic $99.08\%$ $94.81\%$
[…]
real (1.5) $88.79\%$ $61.11\%$
real (ResNet) $95.59\%$ $66.23\%$
real (Opt) $98.10\%$ $68.30\%$
However, we find many terminal blocks to be misclassified as their smaller-sized variant, e.g., PT 4 as PT 2,5 or PT 4-TWIN as PT 2,5-TWIN, and that the number of undetected objects is high (see the appendix). Since this observation does not correspond to the results of the synthetic test and the real-world dataset tends to show the terminal strips on a smaller scale than the synthetic dataset, we assume that the size of the terminal strip components in the image is crucial to the distinction of similar objects that mainly differ in size. Hence, we briefly test this hypothesis by resizing all real images with a constant scaling factor of 1.5 and obtain a performance increase of more than 13% $mAP@0.5$ for both object detectors, as expected. Nevertheless, the sim-to-real performance gap is still significant in this experiment, and a more dynamic scaling approach would be required in practice anyway.
One possibility for dynamically determining the scaling factor for a terminal strip image is to use a preprocessing DL model to predict it. Therefore, we again rendered synthetic images to train a ResNet50 regression model, accepting the additional influence of the domain gap on the detection performance. Utilizing this model to resize the real test images further improves the detection performances of both detectors significantly, reaching 90.75% and 95.59% $mAP@0.5$. This quantifies the domain gap to 8.51% for RetinaNet and 3.49% for Faster R-CNN, which is a remarkable result.
However, the potential of scaling factors has not yet been fully exploited, as demonstrated by the last experiment. Concretely, the mean absolute errors between the optimized and the predicted scaling factors are 0.11 for RetinaNet and 0.13 for Faster R-CNN, preventing the models from another 5.82% and 2.51% $mAP@0.5$ leap in performance. Thus, with optimized image resizing the sim-to-real domain gap reduces to 2.69% and 0.98% $mAP@0.5$ demonstrating the power of combining DR with domain knowledge even in challenging industrial applications.
Figure 6: Example prediction of Faster R-CNN under optimized scaling conditions showing inaccurate bounding boxes caused by the low contrast between markings and background.
The main performance differences between synthetic and real test images seem to lie in the accuracy of the predicted bounding boxes, as the $mAP@\left[0.50:0.95\right]$ scores of RetinaNet and Faster R-CNN are more than 20% worse in comparison. Looking at the predictions, we observe that the white background often causes inaccurate bounding boxes when a terminal block is equipped with a marking at its upper or lower edge so that there is a low contrast between the object and background, as depicted in Figure 6. Furthermore, we belief that the manual annotations are not as exact as the calculated ones, introducing additional inaccuracies. Hence, the lower performance in $mAP@\left[0.50:0.95\right]$ is explainable and could be addressed in real-world applications, e.g., by ensuring backgrounds deviating in color.
5 DISCUSSION
The results show that a real-world industrial object detection task can be learned from synthetic training images based on already available CAD data. Although the Faster R-CNN seems more suitable for sim-to-real generalization, as it outperforms the RetinaNet in every experiment, the underlying Inception-ResNet consists of more than twice as many parameters as the ResNet50 backbone of RetinaNet, increasing the required hardware and computation time. However, both models perform well under optimized scaling conditions, qualifying this approach for potential use in industrial manufacturing. Additionally, we proved that suitable scaling factors can also be learned from simulated data and expect future work to further improve the image preprocessing.
Regardless, it depends on the requirements of the application if a detection rate of 98.10% $mAP@0.5$ is acceptable. For identifying already mounted terminal strip components in an electrical cabinet, e.g., this performance might be sufficient, whereas stricter regulations apply in quality control. In the latter case, however, simplified conditions such as a fixed camera position and knowledge about the background can be exploited, which we hypothesize has another positive impact on the performance, potentially satisfying these regulations. Hence, industrial requirements seem to be within the realm of possibility, although minor uncertainties of the expected performance remain.
Nevertheless, removing the intractability of collecting and annotating data facilitates the development of industrial vision systems that can solve more complex tasks than classical image processing methods and are much faster than manual inspections by humans. Even though it is not guaranteed that this approach satisfies the various industrial requirements in a specific use case, DL models could still assist manual processes saving costs and time. Therefore, the presented results should be considered as a promising baseline with room for improvements to encourage industrial manufacturers to contemplate image synthesis methods in their applications.
6 CONCLUSIONS
In this paper, we combined domain randomization and domain knowledge in an image synthesis pipeline to investigate the impact of synthetic training data in the industrial application of terminal strip object detection. Therefore, we trained the standard object detectors RetinaNet and Faster R-CNN on synthetic training images and evaluated their performance on a real-world test dataset. The results demonstrated the importance of scaling, specifically when some object classes only differ in size. Regardless, we showed that state-of-the-art image synthesis methods enable sim-to-real generalization with a manageable domain gap even in more complex industrial applications, removing the intractability of data collection that is the main obstacle for current DL-based vision tasks in the industry. Moreover, we provide a dataset of synthetic and real terminal strip images so other detection models and preprocessing methods can be benchmarked in this use case.
Since the performance of the DL models highly depends on image scaling in our experiments, in future work, we focus on learning a suitable preprocessing, including rotations, to reach the performance observed under optimized scaling conditions in a fully predictive manner. Moreover, additional object classes should be considered, as product portfolios of industrial manufacturers often consist of more than 40 components, to assess how the performance changes with the number of classes. Finally, we plan to investigate orthographic images of terminal strips since perspective images complicate the determination of accurate bounding boxes in the edge regions.
APPENDIX
Randomly selected synthetic training images generated by the proposed image synthesis pipeline.
Randomly selected images from the real-world dataset.
Confusion matrix of the Faster R-CNN detection model evaluated on the unscaled real images at IoU and score thresholds of 0.5. While many correct detections can be seen on the diagonal, certain misclassifications often occur. Among them are terminal blocks such as PT 4 or UT 6 that are repeatedly misclassified as their smaller-sized variants. Although we expected these types of false predictions, they could not be observed in the synthetic tests, indicating a sim-to-real discrepancy in scaling.